ID |
Date |
Author |
Type |
Category |
Subject |
15640
|
Fri Oct 23 09:03:43 2020 |
anchal | Update | Electronics | HV coil driver packaged into 2U chassis |
Andrew made a battery-powered 0.7 nVrtHz input-referred noise pre-amplifier for gain of 200. That might help you.
Quote: |
we'd need a preamp with better than 1nV/rtHz to directly measure the noise I guess.
|
RXA: 0.7 nV is OK if you're not interested in low noise measurements. Otherwise, we have the transformer coupled pre-amp from SRS which does 0.15 nV/rHz and the Rai Weiss FET amp which has 0.35 nV for high impedance sources. |
15639
|
Thu Oct 22 22:01:53 2020 |
gautam | Update | Electronics | HV coil driver packaged into 2U chassis |
It's not so easy to directly measure this I think, because the filtering is rather aggressive. Attachment #1 shows the measured transfer function (dots) vs the model and Attachment #2 shows the noise. I think this checks out - but I can't definitively rule out some excess noise at 100 Hz from this stage. Because the gain of the HV stage is x31, we'd need a preamp with better than 1nV/rtHz to directly measure the noise I guess. The Acromag noise model in Attachment #2 is based on a measurement I describe here.
Quote: |
what is the noise level before the HV stage? i.e. how well is the acromag noise being filtered?
|
|
Attachment 1: DACnoiseFilterGain.pdf
|
|
Attachment 2: DACnoiseFilterNoises.pdf
|
|
15638
|
Thu Oct 22 13:04:42 2020 |
rana | Update | Electronics | HV coil driver packaged into 2U chassis |
what is the noise level before the HV stage? i.e. how well is the acromag noise being filtered? |
15637
|
Thu Oct 22 11:48:08 2020 |
Yehonathan | Update | BHD | Monte Carlo Simulations |
I found this H1 alog entry by Izumi confirming that the calibrated channels CAL-CS_* need the same dewhitening filter.
This encouraged me to download the PRCL and MICH data and using Jon's example notebook. I incorporated these noise spectra into the MCMC simulation. The most recent results are attached.
I am still missing:
- Laser frequency noise
- Laser RIN
- Estimation of the LO phase noise
- Estimation of the BHD breadboard angular noise
Also, now the MCMC repeats a simulation if it doesn't pass the RF PDs test so the number of valid simulations stays the same. I'm still not sure about why the A+ simulations are much more robust to these tests than aLigo simulations. |
Attachment 1: MICH_AplusMCMC.pdf
|
|
Attachment 2: PRCL_AplusMCMC.pdf
|
|
Attachment 3: SRCL_AplusMCMC.pdf
|
|
Attachment 4: OMC_Comm_AplusMCMC.pdf
|
|
Attachment 5: OMC_Diff_AplusMCMC.pdf
|
|
Attachment 6: OMC_Angle_Yaw_AplusMCMC.pdf
|
|
Attachment 7: OMC_Angle_Pitch_AplusMCMC.pdf
|
|
Attachment 8: Main_Laser_RIN_AplusMCMC.pdf
|
|
15636
|
Thu Oct 22 11:14:47 2020 |
gautam | Update | Electronics | HV coil driver packaged into 2U chassis |
I packaged the HV coil driver into a 2U chassis, hoping for better shielding from pickup. There is still considerable excess noise in measurement vs model around 100 Hz, see Attachment #1. The projected displacement noise from this noise contribution is shown in Attachment #2 - I've also plotted the contribution from the 4.5kohm (planned value for fast path series resistance) for comparison. Attachment #3 has some photos of the measurement setup so if someone sees some red flags, please let me know.
- The noise was measured with the output load connected to a 20ohm load resistor, to simulate an OSEM.
- The input signal was driven with an Acromag, to try and mimic the actual operating conditions as closely as possible (although the fast path input was left unconnected).
- The KEPCO switching HV power supplies were used to power the unit.
I've run out of ideas to try and make the measurement cleaner - the presence of the rather prominent power line harmonics suggests that this is still not perfect, but what more shielding can we implement? I have to make the measurement on the circuit side of the 25 kohm series resistor, so I am using some Pomona minigrabbers to clip onto the leg of the wirewound resistor (see photos in Attachment #3), so that's not great maybe, but what's the alternative?
So if this is truly the noise of the circuit, then while it's an improvement on the current situaiton, it's unsatisfying that such a simple circuit can't match the design expectations. But how do we want to proceed? |
Attachment 1: HVampNoise_driven_chassis.pdf
|
|
Attachment 2: HVampNoise_dispUnits.pdf
|
|
Attachment 3: D1900163_measurementSetup.zip
|
15635
|
Tue Oct 20 20:12:18 2020 |
Koji | Summary | General | DJI OSMO Pocket Camera Kit |
I set up an action cam (DJI OSMO Pocket) and brought it back to the 40m. The kit is now placed in the control room cabinet together with the Canon DSLR.
I might have left the USBC chaging cable at home this time. Will bring it back next time.-> The cable was returned to the kit on Oct 23rd. |
Attachment 1: 20201020200929_IMG_0173.JPG
|
|
15634
|
Mon Oct 19 15:40:02 2020 |
Koji | Update | PEM | Alaska EQ M7.5 |
Alaska M7.5 20:54UTC https://earthquake.usgs.gov/earthquakes/eventpage/us6000c9hg/executive
I looked at the suspensions. The watchdogs have not been tripped.
IMC was locked but continually shaken. (and occasional unlock) |
15633
|
Mon Oct 19 15:38:42 2020 |
Koji | Update | Electronics | Loan: A file binder "40m wiring diagram" |
I'll bring a file binder "40m wiring diagram" to home at the next chance.
There is another one on the shelf in the control room.
(I thought I put it in my bag, but it looks like that I left it somewhere around the fax area) |
15632
|
Fri Oct 16 19:44:41 2020 |
anchal | Summary | General | Lab Entry Notification |
I entered 40m today at around 1:10 pm and left by 1:50 pm. I entered 104 through the machine shop entry. I took top view single picture photos of ITMY, BS, AP, ITMX, ETMX and ETMY tables. The latest photos will be put here on the wiki soon. |
15631
|
Fri Oct 16 09:16:37 2020 |
Yehonathan | Update | BHD | Monte Carlo Simulations |
Pushed another update to MCMC simulation. This includes:
- Added new imbalances: ITM transmission, ITM & ETM RoCs.
- Added new static offsets: DHARD, DSOFT, CHARD, CSOFT. All pitch. The RMS is calculated from the data Jon fetched with /input_noises/input_noises.ipynb.
- SRCL noise ASD and RMS are now taken from data in /input_noises.
- RF PD diagnostics were redone: Instead of post-discarding marginal simulations, simulations are now discarded when one or more of the RF PDs demodulated signal does not cross zero when the associated DOFs are scanned by 1um in the offset state.
The DOFs<->RFPD associations I use are:
DARM |
AS_f2_I |
CARM |
REFL_f1_I |
MICH |
POP_f2_Q |
PRCL |
POP_f1_I |
SRCL |
REFL_f2_I |
However, one thing that bothers me is that for some reason ~ 15 out of 160 aLigo simulations are discarded while none for A+. It can also be seen that the A+ simulations are more spread-out which might be related.
The new simulation results are attached. |
Attachment 1: MICH_AplusMCMC.pdf
|
|
Attachment 2: PRCL_AplusMCMC.pdf
|
|
Attachment 3: SRCL_AplusMCMC.pdf
|
|
Attachment 4: OMC_Comm_AplusMCMC.pdf
|
|
Attachment 5: OMC_Diff_AplusMCMC.pdf
|
|
Attachment 6: OMC_Angle_Yaw_AplusMCMC.pdf
|
|
Attachment 7: OMC_Angle_Pitch_AplusMCMC.pdf
|
|
Attachment 8: Main_Laser_RIN_AplusMCMC.pdf
|
|
15630
|
Thu Oct 15 20:00:23 2020 |
Koji | Summary | General | HEPA AC cord replacement |
The AC cord from the PSL HEPA variac to the junction box was replaced.
Now the HEPA is running at 70%
Showed up at the 40m at 7pm
Preparation
- Closed the PSL shutter.
- Closed the innolight shutter
- Turned off the HEPA mains switch
- Checked the HEPA fan rating: 115V 4.5A.
- Brought the thickest power cord from the wall stock: the rating is 125V 15A. This should sufficiently hold two HEPAs.
Cable Replacing
- Rechecked the wire connection. The new cord has green/black/white wires. And the colors agree with the color of the wires in the junction box.
- Removed the existing cord.
- Attached the new cord.
- Checked the variac AC plug. The terminals in the plug look normal and the AC plug looked sufficiently rigid.
- Checked the connection again. = OK
Testing
- Turned on the HEPA mains switch
- VairAC turned to 70%
- Checked the air flow - The HEPA fans are sucking the air = OK
Closing the work
- Closed the junction box.
- Cleaned up the roof
- Opend the innolight shutter
- Opened the PSL shutter
- Locked the PMC
- Locked the IMC - found the transmission was ~80% of the pre-work due to misalignment of the PMC
- Aligned the PMC - this recovered the IMC REFL of ~5.2 when the IMC was unlocked
Leaving the 40m at 9:30pm
Memo: 40m wiring/Mask/Camera/Red Pitaya/Particle Counter |
Attachment 1: P_20201015_200732.jpg
|
|
Attachment 2: P_20201015_200752.jpg
|
|
Attachment 3: P_20201015_202615.jpg
|
|
Attachment 4: P_20201015_204234.jpg
|
|
15629
|
Thu Oct 15 13:48:58 2020 |
anchal | Summary | General | Lab Entry Notification |
I entered 40m today at around 1:20 pm and left by 1:45 pm. I entered 104 through the machine shop entry. I did the following:
- I took photos and videos of the PSL table with lights on.
- I uncovered the AP table, took photos and video, and covered it back.
- I went to the X End table and took a video without opening the enclosure.
- Apart from flipping light switches, nothing else should have changed.
|
15628
|
Thu Oct 15 10:42:39 2020 |
gautam | Update | BHD | More investigation into RF44 sensing |
Summary of discussion between Koji and gautam on 14 Oct:
- Koji questioned the accuracy of the "open loop" ASD shown here. While it may not be entirely accurate to compute the free-running (homodyne) phase noise simply by taking the arctangent of the I and Q signals (because the magnitude of the signal is also changing), gautam claims the estimate is probably still close to the true homodyne phase, especially since the ratio of the "in-loop" and free-running ASDs gives something that closely approximates the magnitude of the supposed OLG of the system.
- Koji suggested the following tests:
- Investigate the relative stability of the two RF signal generators involved in this system. Since the 44 MHz electrical LO signal (for demodulation) is generated by a separate IFR from the one used to imprint 11 MHz and 55 MHz phase modulation sidebands on the main PSL beam, we want to investigate what the drift is.
- Try implementing an analog feedback loop using LB1005 - the idea being we should be able to implement higher bandwidth control, for better suppression of the high frequency noise (which looking at the ASD is not only due to seismic phase modulation of the IFO output field). Maybe some combination of this and the Marconi investigation would suggest why we have these forests of lines in the ASDs of the error signal?
- Turn off the HEPAs on the PSL enclosure completely as a test, to see if that improves (i) phase noise due to air currents and (ii) mechanical pickup on the fiber producing phase noise.
I tried all of these last night / overnight, here are my findings.
Analog locking of the homodyne phase:
See Attachment #1.
- RF44_I was used as the error signal.
- The "C1:OMC-ZETA_IMON_OUT" channel is actually looking at the error signal monitor from the LB1005, and is uncalibrated in this plot.
- The "monitor" port on the demodulator board provides a convenient location for us to route the demodulated signal to an LB1005 box, while simultaneously digitizing both demodulated quadratures.
- Empirically, I found settings that could engage the lock. I also found that I couldn't increase the gain much more without destroying the lock.
- The time domain signals look much "cleaner" in this analog feedback loop than when I achieved similar stabilization using the digital system. But I will quantify this more when I post some spectra of the in loop error signals.
- I will do some more characterization (loop TF measurement, error point spectrum in lock etc), but in summary, it looks like we still only have ~100 Hz UGF. So something in the loop is limiting the bandwidth. What could it be?
- The main problem is that the LB1005 isn't well suited to remote enabling/disabling of the lock, so this isn't such a great system.
Relative stability of two IFR2023s synchronized to the same FS725 Rb standard:
The electrical LO signal for demodulation of the 44 MHz photocurrent is provided by an IFR2023 signal generator. To maintain a fixed phase relation between this signal, and the phase modulation sidebands imprinted on the interferometer light via a separate IFO2023 signal generator, I synchronize both to the same Rb timing standard (a 10 MHz signal from the FS725 is sent to the rear panel frequency standard input on the IFR). We don't have a direct 44 MHz electrical signal available from the main IFO Marconi at the LSC rack (or anywhere else for that matter). So I decided to do this test at 55 MHz.
- RF input of the demodulator was driven by 5*11.066209 MHz pickoff from the LSC rack.
- LO input of the demodulator was driven by 5*11.066209 MHz signal from the IFR2023 used for the RF44 demodulation setup.
- The outputs were monitored overnight. The RF44_Q channel had a DC level of nearly 0. So this channel is nearly a linear sensor of the phase noise between LO and RF signals.
- To convert ADC counts to radians, I offset the LO Marconi frequency by 100 Hz, and saw that the two quadratures showed pk-pk variation of ~24000cts. So, at the zero crossing, the conversion is 1/(24000/2) rad/ct ~83urad/ct.
- The result is shown in Attachment #2. The "measurement noise" trace corresponds to the RF. input of the demodulator being terminated to ground with a 50 ohm terminator.
- For comparison, I also overlay the phase noise estimate of an individual IFR from Rana. In his investigation, the claim is that the PLL that locks the IFR to the Rb timing standard has ~1kHz UGF, but if my measurement is correct, the relative stability between the two signal generators synchronized to the same timing standard already. degrades at ~1 Hz. Could be just a cts/rad calibration error I guess.
- In any case, we are far from saturating this limit in the homodyne phase lock.
- There are several sharp lines in this measurement too - but I don't know what exactly the source is. Of course the two marconis are plugged into separate power strips, so that may explain the 60 Hz lines and harmonics, but what about those that aren't a multiple of 60 Hz?
A look at the time domain signal:
With the Michelson locked on the dark fringe, the RF44 I and Q signals in the time domain are shown in Attachment #3 for a 1 minute stretch.
- The RF44 signal level bottoms out at ~40 cts. Okay, so this is the offset.
- However, the maximum value of the RF44 signal amplitude seems to be modulated in time. How can we explain this?
|
Attachment 1: analogZetaLock.png
|
|
Attachment 2: relPhaseNoise.pdf
|
|
Attachment 3: sigMagPhase.pdf
|
|
15627
|
Wed Oct 14 18:16:27 2020 |
gautam | Update | General | PSL HEPA-->50% |
Per Koji's suggestion, I turned the PSL HEPA Variac to 50% just now, so that the power load through the burnt electrical cable is reduced by 75%. |
15626
|
Wed Oct 14 17:03:55 2020 |
anchal | Summary | ALS | ALS noise budget update - Added whitening filter for ADC |
Koji recommended that I can add whitening filters to suppress ADC noise easily. I added a filter before ADC in ALS loop with 4 zeros at 1.5 Hz and 4 poles at 100 Hz and added a reversed filter in the digital filter of ALS. This did not change the performance of the loop but significantly reduced the contribution of ADC noise above 1 Hz. One can see ALS_controls.yaml for the filter description. Please let me know if this does not make sense or there is something that I have overlooked.
Now, the dominant noise source is DFD noise below 100 Hz and green laser frequency noise above that. For DFD noise, I used data dating back to Kiwamu's paper. The noise contribution from DFD in the model is lower than the latest measured ALS noise budget post on elog. I'll look further into design details and noise of DFD.
Code, data, and schematics |
Attachment 1: ALS_NoiseBudgetUpdate.pdf
|
|
15625
|
Wed Oct 14 13:28:04 2020 |
Koji | Update | COC | ITM/ETM spares in Downs |
The two ITM spares and two ETM spares are together stored in the optic storage (B110) at Downs. c/o Liyuan and GariLynn |
Attachment 1: IMG_3073.jpeg
|
|
15624
|
Tue Oct 13 21:22:29 2020 |
gautam | Update | General | Space cleared in 1Y1 for new FEs |
[JV, GV]
We cleared up some space in the 1Y1 electronics rack to install the 3 new FE machines. I removed the current driver and laser from 1Y1, they are now stored in the E10 cabinet. I will upload some photos to gPhotos soon.
- I think it's good to have all these FEs in one rack (at least the new ones) - we should then hook it up to an ethernet power source, so that we can remotely power cycle them. I think we have long enough cables to interface to expansion chasses / dolphin switches, but if not, I think it's still a good idea to have these machines in 1Y1 as it is the least sensitive area in terms of immunity to bumping some cable during setup work and disturbing the rest of the IFO.
- We found that the rails that the Supermicros shipped with the servers seem to be just a little too narrow - we mounted these in the rack, but had considerable difficulty sliding the server units in. Once they are in, they don't slide smoothly. Is there some special trick to installing these?
- I spent a few minutes trying to get Debian 8 installed on these machines, so that the rest of the setup work could be done remotely - however, there appear to be some firmware issues and so I'm not gonna dive into this.
- I couldn't find a disk image for Debian 8.5 which is what the KT wikl recommends, so the OS I tried to install was Debian 8.11.
- The error that comes up is related to a "stalled CPU" - apparently this is related to some graphics driver issue (there's another forum page that suggests upgrading the BIOS, but I don't think that's the problem here).
- Anyways, this part of the process is only to install some drivers and do the initial setup - these machines will eventually run a diskless boot from the image on FB, so who knows if there will be some other driver issues/hardware-software incompatibilities there 😱 .
- We should also make an effort to set these machines up with IPMI, but I think we first need to install an OS and a CLI to setup the IPMI.
My cursory browsing of the manual suggests that the initial setup maybe can be done without installing an OS, and then subsequent work, including OS install, can be done remotely. If someone reads more in detail and can provide me a step-by-step, I can follow those instructions (if they aren't available to come into the lab). See here for some brief documentation of how to access the IPMI.
|
15623
|
Tue Oct 13 11:13:54 2020 |
gautam | Update | BHD | Investigation into RF44 sensing |
Attachment #1: spectra of the phase noise between LO and IFO output fields sensed using the RF44 signal.
- Measurement setup:
- LO an IFO fields are combined on a beamsplitter, with ~60% mode-matching efficiency.
- One port of the BS goes to a DCPD.
- The other port goes to an RF sensing photodiode, PDA10CF. The spec-ed dark noise NEP is ~12 pW/rtHz at 1.6 um, (so let's say 25 pW/rtHz) and transimpedance is 5kohms into a 50 ohm load. We can convert this to an equivalent sensing noise at the error point of this loop, though it's more likely that the electronics (demod, ADC etc) noise downstream dictate the sensing limit, which I measure by blocking light on the photodiode.
- The demodulation is done on one of the newly received D0902745 boards - this was just a more compact setup than many cascaded minicircuit components. We don't have the hardware to package this into a chassis to shield against electronics noise pickup yet, so I'm using a bench supply to power this for now (via a voltage regulation board, D1000217.
- "Dark Noise" = ASD with no light incident on the photodiode. "LO field only" = ASD with only the LO field incident on the photodiode.
- The "Dark noise" trace and "LO field only" traces are converted from cts/rtHz to rad/rtHz by noting that when the Michelson is locked on a dark fringe, the demodulated RF44 quadratures have a pk-pk amplitude of ~160 cts (corresponding to pi radians of phase shift). Since in these conditions the demodulated quadratures do not undergo any fringe wrapping, I converted the spectra by simple multiplication.
- For the "RF44 open loop" trace:
- The DC offset in the demodulated signal (due to the RF44 signal from the LO field only) is digitally compensated, so that the fringing has (roughly) zero offset.
- The Michelson was locked on a dark fringe, and the demodulated RF44 quadratures were monitored for ~5 mins. Then arctangent (specifically, arctan2 to get the correct quadrant in the IQ plane) of the two signals was taken to convert the fringing signals to phase noise.
Closing a feedback loop:
- Since it seems like we are sensing a signal (below ~1kHz at least), I tried to close a feedback loop (modelled loop shape shown in Attachment #2, it's just a model because I have to guess what the sensing and actuation gains are, and they're both assumed to be flat, digital delays etc aren't accounted for). I've also added the inferred loop gain by taking the ratio of the in loop and unsuppressed ASDs (though of course I don't account for the flat sensing noise at higher frequencies). At least qualitatively, things line up...
- While I can get the light level on the DCPD to stabilitze somewhat, the loop is not at all stable, and the suppression isn't very good at all.
- Not sure how meaningful any of the spectra with the loop closed are, but FWIW, I've put in the spectra of the demodulated RF44 signals with the loop engaged (RF44 Q is used as the error signal). A clear problem is evident at ~120 Hz, and the forest of lines isn't helping for sure. Also unclear to me why the I and Q signals don't have the same profile at low frequencies.
Conclusions/Questions:
- What is the reason for the huge forests of lines in the "RF44 open loop" ASD, that are absent in the other two traces? If this were electrical pickup, it should be there in all three traces?
- Is the shape of the spectrum reasonable? The roll-off above ~5 Hz doesn't seem quite steep enough to be seismic noise from the suspensions. Can it really be that the Michelson dark field has such high phase noise?
- How can we get this scheme to give us cleaner sensing?
- The actuation chain was verified to work fine with the single bounce beam from an ITM interfered with the LO field, and using the DC light level as an error signal and locking to the half-fringe point. So the problem is not due to insufficient actuation range. Seems like the error signal is so polluted with these forests of lines that even though there is some suppression of the error signal at low frequencies, the unsuppressed noise is still significant. I can't solve the problem by simply increasing the loop gain...
- It is not shown here, but with only the LO field incident on the RFPD, I see a drift of the demodulated signals on the ~5 minute timescale - is this just due to fiber length change? If so, this is potentially problematic, as on long time scales, the true zero of the error point of the servo would be changing on the ~5 minute timescale. This would be true even for the final suspended scheme - if the path length between PR2 and the homodyne BS changes by some microns, we would have to correct this at DC?
|
Attachment 1: phaseNoisePSD.pdf
|
|
Attachment 2: loopTF.pdf
|
|
15622
|
Fri Oct 9 18:32:14 2020 |
anchal | Summary | ALS | ALS noise budget update - Updated AUX PDH Loop values |
The only two PZT Phase modulation transfer function measurements I could find are 40m/15206 and 40m/12077. Both these measurements were made to find a good modulation frequency and do not go below 50 kHz. So I don't think these will help us. We'll have to do a frequency transfer function measurement at lower frequencies.
I'm still looking for ALS PDH loop measurements to verify the model. I found this 40m/15059 but it is only near the UGF. The UGF measured here though looks very similar to the model prediction. A bit older measurement in 2017 was this 40m/13238 where I assume by ALS OLTF gautum meant the green laser PDH OLTF. It had similar UGF but the model I have has more phase lag, probably because of a 31.5 kHz pole which comes at U7 through the input low pass coupling through R28, C20 and R29 (See D1400293)
If the green laser is not being used, can I go and take some of these measurements myself? |
15621
|
Thu Oct 8 18:40:42 2020 |
Koji | Update | Computer Scripts / Programs | Finesse GUI |
Is it better than Luxor? https://labcit.ligo.caltech.edu/~jharms/luxor.html |
15620
|
Thu Oct 8 15:08:25 2020 |
gautam | Update | General | Some boxes moved from 40m entry hallway |
- UPS batteries
- 2x HEPA filters
- VWR chemicals (methanol)
These boxes were moved from the 40m hallway to the inside of the VEA so that we have some space to walk around. You can find some pictures here. |
15619
|
Thu Oct 8 11:59:52 2020 |
rana | Summary | ALS | ALS noise budget update - Updated AUX PDH Loop values |
For all the loops where we drive the NPRO PZT, there is some notch/resonance feature due to the PZT mechanical resonance. In the IMC loop this limits the PZT/EOM crossove to be less than 25 kHz. I don't have a model for this, btu it should be included.
If you hunt through the elogs, people have measured the TF of ALS NPRO PZT to phase/frequency. Probably there's also a measured ALS PDH loop somewhere that you could use to verify your model. |
15618
|
Thu Oct 8 08:37:15 2020 |
gautam | Update | Computer Scripts / Programs | Finesse GUI |
This looks cool, we should have something similar, can be really useful. |
15617
|
Wed Oct 7 16:56:23 2020 |
anchal | Summary | ALS | ALS noise budget update - Updated AUX PDH Loop values |
AUX PDH Loop update
I used D1400293 to get the latest logged details about the universal PDH box used to lock the green laser at X end. The uPDH_X_boost.fil file present there was used to obtain the control model for this box. See attachment one for the code used. Since there is a variable gain stage in the box, I tuned the gain of the filter model F_AUX in ALS_controls.yml to get the maximum phase margin in the PDH lock of the green laser. Unity gain frequency of 8.3 kHz can be achieved in this loop and as Gautam pointed out earlier, it can't be increased much further without changes in the box.
ALS Noise Budget update
The ALS control model remains stable with a reduction in total estimate noise because of the above update. There are few things to change though:
- This model is for a single arm locking where the beatnote signal between green laser and frequency doubled main laser is fed back to ETM at X end. Currently, Gautam is using a different scheme to lock where the feedback is sent to PSL-MC loop and the beat is taken between IR signals.
- In the LSC controls, I couldn't find a place where the digital ALS filter I have been optimizing and Kiwamu used, was placed. From what I gathered, after demodulation of beat note signal, a digital PLL is employed and the error signal is few to the Servo Filters directly. I might be missing some script which specifically switches on a particular set of filter modules in the XARM/YARM path when arms are locked through ALS.
- Another straight forward job for me is to verify the PSL-MC loop parameters with he TTFSS used. I'll do this next.
|
Attachment 1: Extract_X_AUX_PDH_Model.zip
|
Attachment 2: ALS_NoiseBudgetUpdate.pdf
|
|
15616
|
Wed Oct 7 13:06:27 2020 |
Koji | Update | General | Presence in the lab |
Tue evening from 4pm~6pm, Koji made a social distant tour for Anchal. We were present around the PSL/AS/ETMX tables. |
15615
|
Tue Oct 6 14:35:16 2020 |
Jordan | Update | VAC | Spare forepumps |
I have placed 3 new in box, IDP 7 forepumps along the x arm of the interferometer. These are to be used as spares for both the 40m and Clean and Bake. |
15614
|
Tue Oct 6 07:37:20 2020 |
yehonathan | Update | Wiki | New TIS measurements of 40m Optics |
LiYuan has kindly done some Total Integrating Sphere (TIS) measurements on ITMU01 and ITMU02. A summary of the measurement is attached. I uploaded the measurements and some analysis script to nodus at /home/export/home/40m_TIS. I created a Wiki page for the measurements and linked to it from the core optics page.
These TIS measurements look very similar to the TIS of the LIGO optics. Further analysis shows that the scatter loss is 10+/-1.7 ppm for ITMU01 and 8.6+/-0.4 ppm for ITMU02.
In this calculation, a gaussian beam the same size bouncing off the 40m ITMs is assumed to scatter from the mirrors. The error is calculated by moving the beam around randomly with STD of 1mm.
In LiYuan's setup, TIS is measured for scattering angles between 1 and 75 degrees. If we go further and assume that the scatter is Lambertian we can extrapolate that the total loss is 10.9+/-1.9 ppm for ITMU01 and 9.2+/-0.5 ppm for ITMU02.
These measurements complete the loss budget nicely since with the 6ppm loss predicted from the phase maps, the total loss in the arm cavities would be 6+10+10=26ppm which is very close to the 28ppm loss that was measured after the arm cavity optics were cleaned. |
Attachment 1: ITMU_sn01-02_tis_1.pdf
|
|
15613
|
Mon Oct 5 14:01:41 2020 |
gautam | Update | Electronics | aLIGO demod boards stuffed and delivered |
We received 20pcs of stuffed demodulator boards from Screaming Circuits today. Some caveats:
- The AP1053 amplifiers weren't stuffed. Note that this part is no longer in standard production, and lead time for a custom run is ~half a year. I recommend stuffing R2 and using a minicircuits amplifier upstream of this board. We have 6 pcs of AP1053 in hand so we can use those for the first AS WFS, but a second WFS will require some workaround.
- AD8306ARZ weren't sent to Screaming Circuits. This part is used for the LO and RF signal level detection/monitoring stage, and so aren't crucial to the demodulation operation. @Chub, did we order the correct part now? They are rather pricey so maybe we can just adapt the footprint using some adaptor board?
- DQS-10-100 hybrid 90 degree splitters were delivered to us after the lot was sent to Screaming Circuits. We have the pieces in hand, so we can just stuff them as necessary.
I removed 1 from the group to stuff some components that weren't sent to Screaming Circuits and test the functionality on the benchtop, the remaining have been stored in a plastic box for now as shown in Attachment #1. The box has been delivered to Chub who will stuff the remaining 19 boards once I've tested the one piece. |
Attachment 1: IMG_8888.JPG
|
|
15612
|
Mon Oct 5 00:53:16 2020 |
Koji | Update | BHD | Single bounce interferometer locked |
🤘🤘🤘
|
15611
|
Mon Oct 5 00:37:19 2020 |
gautam | Update | BHD | Single bounce interferometer locked |
Summary:
The simple interferometer, composed of a single bounce reflection from ITMY and the LO beam deilvered via fiber to the AS table, can be locked - i.e. the phase of the LO beam can be controlled such that the DC light level on the DCPDs after the two beams are interfered can be stabilized. This test allows us to confirm that various parts of the sensing and actuation chain (e.g. PI PZT for homodyne phase control, Trek amplifier etc etc) are working.
I will post more quantitative analysis tomorrow.
Optical configuration:
- LO beam is a pickoff of the main PSL beam from just before it goes into the vacuum. The optical power arriving on each DCPD after the various beamsplitters, coupling loss etc is ~200 uW.
- IFO beam is the single bounce reflection from ITMY. For this test, ETMY, ITMX and ETMY are misaligned. Optical power arriving on each DCPD is ~80uW.
- The two beams are interfered on a 50-50 beamsplitter. The mode-matching efficiency was estimated to be ~50% which isn't stellar, but should be fine for this test.
- So, at half-fringe, we expect the signal on each DCPD to be linearly proportional to the phase difference between the two fields, and so we can use that as an error signal.
Servo topology:
Attachment #2 shows the servo topology.
- For a first attempt to close the feedback loop, we can consider the two blocks labelled "Sensing Chain" and "Actuation chain" to have a flat frequency response. While this isn't true, for a taget loop with ~100 Hz UGF, I think the approximation is reasonable.
- From the peak-to-peak value (160 cts) of the DCPD signals when the homodyne phase is uncontrolled, I estimate a sensing response (at half-fringe) of approximately 0.3 ct/nm, since this corresponds to 532nm of relative phase between the two beams.
- An inverting summing amplifier is used to map the +/- 2^15 ct DAC range to 0-125V on the PI PZT. Assuming the full stroke of the PZT is 10um per the datasheet, and that this voltage range drives half of the full stroke (this is just a guess since all the old PI PZT circuits were designed to work at 0-250 V), we get an actuation coefficient of 0.075 nm/ct.
- Using these two numbers, we can then design a digital feedback loop that gives an open loop transfer function with ~100 Hz UGF, and sufficient stability margin.
- From the earlier measurements, we have an estimate for the amount of phase fluctuations caused by (i) seismic disturbances and (ii) fiber phase noise. This is the quantity we wish to suppress, and the suppression factor will be 1/(1+L), where L is the open loop gain.
- I didn't do this in any systematic way, but the loop in Attachment #3 seemed like a reasonable shape that would suppress the error signal RMS by ~10x, as shown in Attachment #4. So I decided to try this out.
Other notes:
- The idea of offloading the DC control voltage to the ITMY suspension seemed to work fine.
- It also seems like the relative phase between the two beams doesn't drift by so large an amount in short time scales, at least at night/quiet seismic conditions. So it is possible to maintain the lock for several seconds without having to offload the DC signal to the suspensions.
- I didn't bother adapting the FSS Slow PID script to do this offloading in an automated way, seemed like more trouble than was just doing it by hand. But we may want to automate this in the future.
- I couldn't make a clean measurement of the loop transfer function using the usual IN1/IN2 method. Introducing a step offset at the error point, the servo is able to track it (I didn't fit the step response time, but it's not as if the loop bandwidth is <1 Hz or something).
I have to compare the measured in-loop error signal ASD to the free-running one to get a feel for what the UGF is, I guess, to rule out a weird loop.
- Update 1100 Oct 6 2020: I have now added measured, in-loop, error point spectra to Attachment #4. Looks like there might be significant sensing noise re-injection.
- Initially, I forgot to turn the HEPA on the PSL down for the measurement. So I have the two traces to compare. Looks like with the HEPA turned up to full, there is more noise in the 50-200 Hz range.
- The trace marked "highGain" was taken with an overall loop gain that was 3dB higher than the nominal value - I could see some oscillations start to appear, and in the spectrum, maybe the feature at ~150 Hz is evidence of some gain peaking?
Conclusions:
- The PI PZT seems to work just fine.
- Need to look into the loop shape. I guess it's not reasonable to expect a UGF much higher than 100-200 Hz, because of the various delays in the system, but maybe the low frequency suppression can be made better.
- What are the next steps?? What does this mean for the RF44 sensing scheme?
|
Attachment 1: simpleHomodyne.png
|
|
Attachment 2: singleBounceIFO.pdf
|
|
Attachment 3: proposedController.pdf
|
|
Attachment 4: freeRunningSuppressed.pdf
|
|
15610
|
Sun Oct 4 15:32:21 2020 |
gautam | Update | SUS | Suspension health check |
Summary:
After the earthquake on September 19 2020, it looks to me like the only lasting damage to suspensions in vacuum is the ETMY UR magnet being knocked off.
Suspension ringdown tests:
I did the usual suspension kicking/ringdown test:
- One difference is that I now kick the suspension "N" times where N is the number of PSD averages desired.
- After kicking the suspension, it is allowed to ring down with the damping disabled, for ~1100 seconds so that we can get spectra with 1mHz resolution.
- We may want to get more e-folding times in, but since the Qs of the modes are a few hundred, I figured this is long enough.
- I think this kind of approach gives better SNR than letting it ringdown 10,000 seconds (for 10 averages with 10 non overlapping segments of 1000 seconds), and I wanted to test this scheme out, seems to work well.
- Attachment #1 shows a summary of the results.
- Attachment #2 has more plots (e.g. transfer function from UL to all other coils), in case anyone is interested in more forensics. The data files are large but if anyone is interested in the times that the suspension was kicked, you can extract it from here.
Conclusions:
- My cursory scans of the analysis don't throw up any red flags (apart from the known problem of ETMY UR being dislodged) 👌 .
- The PRM data is weird
- I believe this is because the DC bias voltage to the coils was significantly off from what it normally is when the PRC is aligned.
- In any case, I am able to lock the PRC, so I think the PRM magnets are fine.
The PRC angular FF no longer works turns out this was just a weird interaction with the Oplev loop because the beam was significantly off-centered on the Oplev QPD. Better alignment fixed it, the FF works as it did before.
With the PRC locked and the carrier resonant (no ETMs), the old feedforward filters significantly degrade the angular stability to the point that the lock is lost.
My best hypothesis is that the earthquake caused a spot shift on PR2/PR3, which changed the TF from seismometer signal to PRC spot motion.
Anyways, we can retrain the filter.
- The fact that the PRC can be locked suggest PR2/PR3 are still suspended and okay.
- The SRM data is also questionable, because the DC bias voltage wasn't set to the values for an aligned SRC when the data was collected
- Nevertheless, the time series shows a clean ringdown, so at least all 5 OSEMs are seeing a signal.
- Fact that the beam comes out at the AS port suggest SR3/SR2 suspensions are fine 👍
Attachment #2 also includes info about the matrix diagonalization, and the condition numbers of the resulting matrices are as large as ~30 for some suspensions, but I think this isn't a new feature. |
Attachment 1: combined.pdf
|
|
Attachment 2: allPlots.zip
|
15609
|
Sat Oct 3 16:51:27 2020 |
gautam | Update | CDS | RFM errors |
Attachment #1 shows that the c1rfm model isn't able to receive any signals from the front end machines at EX and EY. Attachment #2 shows that the problem appears to have started at ~430am today morning - I certainly wasn't doing anything with the IFO at that time.
I don't know what kind of error this is - what does it mean that the receiving model shows errors but the sender shows no errors? It is not a new kind of error, and the solution in the past has been a series of model reboots, but it'd be nice if we could fix such issues because it eats up a lot of time to reboot all the vertex machines. There is no diagnostic information available in all the places I looked. I'll ask the CDS group for help, but I'm not sure if they'll have anything useful since this RFM technology has been retired at the sites (?).
In the meantime, arm cavity locking in the usual way isn't possible since we don't have the trigger signals from the arm cavity transmission.
Update 1500 4 Oct: soft reboots of models didn't do the trick so I had to resort to hard reboots of all FEs/expansion chassis. Now the signals seem to be okay. |
Attachment 1: RFMstat.png
|
|
Attachment 2: RFMerrs.png
|
|
15608
|
Fri Oct 2 12:52:22 2020 |
gautam | Update | General | Spectroscopic grade Isopropanol delivered |
2x500 ml bottles of spectroscopic grade isopropanol were delivered. I marked them with today's date and placed them in the solvent cabinet. In the process, my shoulder bumped the laser interlock switch by the door to the VEA in the drill press area, which turned the PSL NPRO off. I turned it back on just now. The other NPROs are not connected to the interlock and so were unaffected. |
15607
|
Fri Oct 2 10:29:49 2020 |
gautam | Update | Optical Levers | ETMY, BS and ITMX HeNes degrading |
Attachment #1 shows that the ITMX, ETMY and beamsplitter Oplev light levels have decayed significantly from their values when installed. In particular, the ETMY and ITMX sum channels are now only 50% of the values when a new HeNe was installed. ELOG search revealed that ITMY and ETMX HeNes were replaced with newly acquired units in March and September of last year respectively. The ITMX oplev was also replaced in March 2019, but the replacement was a unit that was being used to illuminate our tourist attraction glass fiber at EX.
We should replace these before any vent as they are a useful diagnostic for the DC alignement reference. |
Attachment 1: OLsum.png
|
|
15606
|
Thu Oct 1 17:44:48 2020 |
gautam | Update | General | Some inventory notes |
The optomechanics stock in the lab was in a sad state. We have obtained the following from Thorlabs in the last two months:
- 6 pcs each of DT12 and DT12B compact translation stages (for lens mode matching).
- 3pcs each KM100PM + PM3 cube beamsplitter mounts (for polarization control).
- 1 Post / spacer kit for height adjustment.
- 3 pcs ea of K6XS + AD11F + F220APC for fiber applications.
I have used some of these for the ari BHD setup. The unused items are stored in the shelves that house the optomechanics ~halfway down the east arm. I'm wondering what's a good setup to document the stock of this stuff so we can always have a healthy stock of optomechanics (at least the non speciality ones like posts, spacers etc). It sucks to realize at 0000hrs that you're missing a 3mm shim or 250mm converging lens or something. |
15605
|
Wed Sep 30 19:45:56 2020 |
rana | Update | General | HEPA blower startup capacitor replacement |
it would be a good idea for us to have an auto-reminder to have us check the flow of all the HEPAs in the lab and elog it once a year so that we can replace filters and pre-filters appropriately. |
15604
|
Wed Sep 30 17:12:24 2020 |
gautam | Update | General | HEPA blower startup capacitor replacement |
[JV, GV]
The HEPAs work again. After running the HEPAs for ~1 hour, I checked the particle count on the PSL table - the meter registered 0 for both 0.3 um and 0.5 um. So I decided to turn the NPRO back on, at ~1730 local time. The PMC and IMC were readily locke, so the basic interferometer functionality is returned, and we can now go ahead with (i) vent prep (ii) air BHD tests and (iii) IMC debuggin as was discussed on the call today. The earth is shaking, but nothing serious so far, I will resume alignment of the interferometer later in the evening when hopefully things have calmed down a bit more...
Procedure:
- Turned off mains switch on the NW corner of the PSL enclosure. Then, disconnected the power cables from mains to Variac and from Variac to HEPAs. Made sure both HEPAs were set to "OFF".
- With confidence that no AC power was reaching the motors, I removed the pre-filters, and removed the old startup capacitors. These don't have a polarity, but I marked the cable that was connected to the left terminal of the capacitor when the cap is viewed with the label facing you, in the interest of changing as few things as possible.
- The two new capacitors were measured with the LCR meter - the meter registered ~7.5 uF, as expected. Unsurprisingly, the old capacitors that were removed didn't register any reading on the LCR meter. The terminals weren't shorted, but I don't know what the failure mode for this kind of capacitor is.
- The two new capacitors were installed. Then, I tested the system by undoing all the changes in bullet #1. We found that the Variac needs to be set to 100% for the motors to startup.
- The motor speed was found to vary as the Variac dial was turned. FWIW, at the "nominal" setting of 33% on the Variac (when we run the interferometer), I could see both fan blades were turning, but the flow was low enough that you couldn't hear any wind (at least, neither Jordan nor I could).
- Turned off the mains agian, and cleaned up - restored the insulating rubber sleeve on the capacitor leads, and re-installed the pre-filters on the HEPA blowers. Then we turned both blowers back on.
Note that the many other issues Koji noted in the preceeding elog (e.g. flaky wiring) have not been addressed.
Flow measurements:
Chub kindly provided us with an electronic anemometer. With the meter held directly against the HEPA filter inside the enclosure, we measured ~700 cfm of airflow on each of the two HEPAs, with the Variac set to 100% and the HEPAs themselves set to "High". With the Variac at 50%, the flow drops to ~160 cfm. At the nominal setting of 33%, the meter didn't register any flow. I don't know what the spec'd flow rate is for this combination of blower + filter, but Jordan says similar units in Downs register ~1500 cfm at the "High" setting. The two protable (similarly sized) HEPA units in the 40m, one at ITMY and one at ETMY, register ~900 cfm and ~1100 cfm respectively, when set to high. So we may want to revisit what the "nominal" HEPA setting should be, in case the filters have become clogged over time.
Some photos of the HEPA blowers with the pre-filters off and the capacitors switched out may be found here. |
15603
|
Tue Sep 29 17:07:25 2020 |
gautam | Update | General | Lab visit for inventory location |
I was in the lab from 1630-1830. I have located and visually inspected all the parts required for the magnet regluing / optic cleaning parts of the planned vent, except the fresh batches of scpectroscopic grade solvents. I was in the cleanroom part of the clean and bake lab from 1630-1700. |
15602
|
Wed Sep 23 15:06:54 2020 |
Jordan | Update | VAC | TP2 Forepump Re-install |
I removed the forepump to TP2 this morning after the vacuum failure, and tested in the C&B lab. I pumped down on a small volume 10 times, with no issue. The ultimate pressure was ~30 mtorr.
I re-installed the forepump in the afternoon, and restarted TP2, leaving V4 closed. This will run overnight to test, while TP3 backs TP1.
In order to open V1, with TP3 backing TP1, the interlock system had to be reset since it is expecting TP2 as a backing pump. TP2 is running normally, and pumping of the main volume has resumed.
gautam 2030:
- The monitor (LCD display) at the vacuum rack doesn't work - this has been the case since Monday at least. I usually use my laptop to ssh in so I didn't notice it so it could have been busted from before. But for anyone wishing to use the workstation arrangement at 1X8, this is not great. Today, we borrowed the vertex laptop to ssh in, the vertex laptop has since been returned to its nominal location.
- The modification to the interlock condition was made by simply commenting out the line requiring V4 to be open for V1 to be opened. I made a copy of the original .yaml file which we can revert to once we go back to the normal config.
- I also opened VM1 to allow the RGA scans to continue to be meaningful.
- At the time of writing, all systems seem nominal. See Attachment #2. The vertical line indicates when we started pumping on the main volume again earlier today, with TP3 backing TP1.
Unclear why the TP2 foreline pump failed in the first place, it has been running fine for several hours now (although TP2 has no load, since V4 isolates it from the main volume). Koji's plots show that the TP2 foreline pressure did not recover even after the interlock tripped and V4 was closed (i.e. the same conditions as TP2 sees right now). |
Attachment 1: Screenshot_from_2020-09-23_15-15-43.png
|
|
Attachment 2: MainVolPumpDown.png
|
|
15601
|
Wed Sep 23 11:13:49 2020 |
anchal | Summary | ALS | ALS noise budget update |
Yes, that loop was unstable. I started using the time domain response to check for the stability of loops now. I have been able to improve the filter slightly with more suppression below 20 Hz but still poor phase margin as before. This removes the lower frequency region bump due to seismic noise. The RMS noise improved only slightly with the bump near UGF still the main contributor to the noise.
For inclusion of real spectra, time delays and the anti-aliasing filters, I still need some more information.
Few questions:
- What anti-aliasing filters are used in ALS?
- Is the digital delay fixed to a constant upper limit or is it left to change as per filters? I have already used a 470 us delay (modeled with Pade 4th order approximation).
- I could not find a good place where channel names are listed with corresponding meaning. Where can I find them?
- Is there a channel which keeps record of lock status? In short, how do I find the in-lock times
Additional Info:
The code used to calculate the transfer functions and plot them is in the repo 40m/ALS/noiseBudget
Related Elog post with more details: 40m/15587 |
Attachment 1: ALS_NoiseBudgetUpdate.pdf
|
|
Attachment 2: ALS_controls.yaml
|
# -----------------------------------------------------------------------------
# AUX
# -----------------------------------------------------------------------------
## Cavity Pole
C_AUX:
p: 1.8883e+04
k: 1.1865e+05
H_AUX:
z: 0
... 113 more lines ...
|
15600
|
Wed Sep 23 10:06:52 2020 |
Koji | Update | VAC | TP2 running HOT |
Here is the timeline. This suggests TP2 backing RP failure.
1st line: TP2 foreline pressure went up. Accordingly TP2 P, current, voltage, and temp went up. TP2 rotation went down.
2nd line: TP2 temp triggered the interlock. TP2 foreline pressure was still high (10torr) so TP2 struggled and was running at 1 torr.
3rd line: Gautam's operation. TP2 was isolated and stopped.
Between the 1st line and 2nd line, TP2 pressue (=TP1 foreline pressure) went up to 1torr. This made TP1 current increased from 0.55A to 0.68A (not shown in the plot), but TP1 rotation was not affected. |
Attachment 1: Screen_Shot_2020-09-23_at_10.00.43.png
|
|
15599
|
Wed Sep 23 08:57:18 2020 |
gautam | Update | VAC | TP2 running HOT |
The interlocks tripped at ~630am local time. Jordan reported that TP2 was supposedly running at 52 C (!).
V1 was already closed, but TP2 was still running. With him standing by the rack, I remotely exectued the following sequence:
- VM1 closed (isolates RGA volume).
- VA6 closed (isolates annuli from being pumped).
- V7 opened (TP3 now backs TP1, temporarily, until I'm in the lab to check things out further).
- TP2 turned off.
Jordan confirmed (by hand) that TP2 was indeed hot and this is not just some serial readback issue. I'll do the forensics later. |
Attachment 1: Screen_Shot_2020-09-23_at_8.55.39_AM.png
|
|
15598
|
Tue Sep 22 23:17:51 2020 |
Koji | Update | General | HEPA Inspection |
Dimensions / Specs
- HEPA unit dimentions
- HEPA unit manufacturer
- Motor
- Capacitor |
Attachment 1: A_HEPA_Dimention.JPG
|
|
Attachment 2: B_HEPA_Company.JPG
|
|
Attachment 3: C_North_HEPA_Spec.JPG
|
|
Attachment 4: D_South_HEPA_Spec.JPG
|
|
Attachment 5: E_Motor_Spec.JPG
|
|
Attachment 6: F_Cap_Spec.JPG
|
|
15597
|
Tue Sep 22 23:16:54 2020 |
Koji | Update | General | HEPA Inspection |
Gautam reported that the PSL HEPA stopped running (ELOG 15592). So I came in today and started troubleshooting.
It looks like that the AC power reaches the motors. However, both motors do not run. It looks like the problem exists in the capacitors, the motors, or both.
Parts specs can be found in the next ELOG.
Attachment 1 is the connection diagram of the HEPA. The AC power is distributed by the breaker panel. The PSL HEPA is assigned to use M22 breaker (Attachment 2). I checked the breaker switch and it was (and is) ON. The power goes to the junction box above the enclosure (Attachment 3). A couple of wires goes to the HEPA switch (right above the enclosure light switch) and the output goes to the variac. The inside of the junction box looked like this (Attachment 4).
By the way, the wires were just twisted and screwed into a metal threaded (but isolated) caps (Attachment 5). Is this legit? Shouldn't we use stronger crimping? Anyway, there was nothing wrong with the caps w.r.t the connection for now.
I could easily trace the power up to the variac. The variac output was just fine (Attachment 6). The cord goes from the variac to the junction box (and then HEPAs) looked scorched. The connection from the plug to HEPAs was still OK, but this should be eventually replaced. Right now the cable was unplugged after the following tests for the safety reason.
The junction box for each HEPA unit was opened to check the voltage. The supply voltage came to the junction boxes and it was just fine. In Attachments 8 & 9, the voltages look low but this is because I just turned the variac only a little.
At the (main) junction box, the resistances of the HEPAs were checked with the Fluke. As the HEPA units are connected to the AC in parallel, the resistances were individually checked as follows.
South HEPA SW |
North HEPA SW |
Resistance |
OFF |
OFF |
High |
OFF |
LO |
5 Ohm |
OFF |
HIGH |
7 Ohm |
LO |
OFF |
7 Ohm |
HIGH |
OFF |
5 Ohm |
The coils were not disconnected (... I wonder if the wiring of South HEPA was flipped? But this is not the main issue right now.)
By removing the pre-filters, the motors were inspected Attachments 10 & 11. At least the north HEPA motor was warm, indicating there was some current before. A capacitor was connected per motor. When the variac was tuned up a bit, one side of the capacitor could see the voltage. I could not judge which has the issue between the capacitor and the motor. |
Attachment 1: 0_PSL_HEPA.pdf
|
|
Attachment 2: 1_Breaker_Panel.JPG
|
|
Attachment 3: 2_Junction_Box.JPG
|
|
Attachment 4: 3_Junction_Box_Inside.JPG
|
|
Attachment 5: 4_Junction_Box_Inside_2.JPG
|
|
Attachment 6: 5_Variac_100%.JPG
|
|
Attachment 7: 6_VariAC_to_HEPA.JPG
|
|
Attachment 8: 7_North_HEPA_IN.JPG
|
|
Attachment 9: 8_South_HEPA_IN.JPG
|
|
Attachment 10: 9_North_Prefilter_Removed.JPG
|
|
Attachment 11: 10_South_Prefilter_Removed.JPG
|
|
15596
|
Tue Sep 22 22:38:11 2020 |
gautam | Update | BHD | Sensing scheme for homodyne phase - some analytic calcs |
I got some feedback from Koji who pointed out that the phase tracker is not required here. This situation is similar to the phase locking of two lasers together, which we frequently do, except in that case, we usually we offset the absolute frequencies of the two lasers by some RF frequency, and we demodulate the resulting RF beatnote to use as an error signal. We can usually acquire the lock by simply engaging an integrator (ignoring the fact that if we actuate on the laser PZT, which is a frequency actuator, just a proportional feedback will be sufficient because of the phase->frequency conversion), the idea being that the error signal is frequently going through a zero-crossing (around which the sinusoidal error signal is approximately linear) and we can just "catch" one of these zero crossings, provided we don't run of actuation range.
So the question here becomes, is the RF44 signal a suitable error signal such that we can close a feedback loop in a similar way? To try and get more insight, I tried to work out the situation analytically. I've attached my thinking as a PDF note. I get some pretty messy complicated expressions for the RF44 signal contributions, so it's likely I've made a mistake (though Mathematica did most of the heavy lifting), it'll benefit from a second set of eyes.
Anyways, I definitely think there is some additional complications than my simple field cartoon from the preceeding elog would imply - the relative phases of the sidebands seem to have an effect, and I still think the lack of the PRC/SRC make the situation different from what Hang/Teng et. al. outlined for the A+ homodyne phase control analysis. Before the HEPA failed, I had tried closing the feedback loop using one quadrature of the demodulated RF44 signal, but had no success with even a simple integrator as the loop (which the experience with the PLL locking says should be sufficient, and pretty easily closed once we see a sinusoidally oscillating demodulated error signal). But maybe I'm overlooking something basic conceptually?
Quote: |
eSummary:
I don't think the proposed scheme for sensing and controlling the homodyne phase will work without some re-thinking of the scheme. I'll try and explain my thinking here and someone can correct me if I've made a fatal flaw in the reasoning somewhere.
|
|
Attachment 1: simpleMich.pdf.zip
|
15595
|
Tue Sep 22 16:29:30 2020 |
Koji | Update | General | Control Room AC setting for continuous running |
I came to the lab. The control room AC was off -> Now it is on.
Here is the setting of the AC meant for continuous running |
Attachment 1: P_20200922_161125.jpg
|
|
15594
|
Tue Sep 22 12:14:42 2020 |
rana | Summary | ALS | ALS noise budget update |
This ALS loop is not stable. Its one of those traps that comes from using only the Bode plot to estimate the loop stability. You have to also look at the time domain response - you can look at my feedback lecture for the SURF students for some functions. |
15593
|
Tue Sep 22 00:14:43 2020 |
anchal | Summary | ALS | ALS noise budget update |
This is not a reply to comments given to the last post; Still working on incorporating those suggestions.
Trying out a better filter from scratch
Rana suggested looking first at what needs to be suppressed and then create a filter suited for the noise from scratch. So I discarded all earlier poles and zeros and just kept the resonant gains in the digital filter. With that, I found that all we need is three poles at 1 Hz and a gain of 8.1e5 gives the lowest RMS noise value I could get.
Now there can be some practical reasons unknown to me because of which this filter is not possible, but I just wanted to put it here as I'll add the actual noise spectra into this model now.
Few questions:
- What anti-aliasing filters are used in ALS?
- Is the digital delay fixed to a constant upper limit or is it left to change as per filters? I have already used a 470 us delay (modeled with Pade 4th order approximation).
- I could not find a good place where channel names are listed with corresponding meaning. Where can I find them?
- Is there a channel which keeps a record of lock status? In short, how do I find the in-lock times
|
Attachment 1: ALS_NoiseBudgetUpdate.pdf
|
|
Attachment 2: ALS_controls.yaml
|
# -----------------------------------------------------------------------------
# AUX
# -----------------------------------------------------------------------------
## Cavity Pole
C_AUX:
p: 1.8883e+04
k: 1.1865e+05
H_AUX:
z: 0
... 106 more lines ...
|
15592
|
Mon Sep 21 16:40:52 2020 |
gautam | Update | PSL | HEPAs are no longer running |
The HEPA filters on the PSL enclosure are no longer running. I tried cycling the power switch on the NW corner of the enclosure, and also turned the dial on the Variac down to zero and back up to maximum, no effect. Judging by the indicator LED on it, the power is making it to the Variac - bypassing the Variac and directly connecting the HEPA to the AC power seems to have no effect either.
I can't be sure, but I'm almost certain I heard them running at max speed an hour ago while I was walking around inside the VEA. Probably any damage that can happen has already been done, but I dialled down the Innolight injection current, and closed its shutter till the issue can be resolved. |
15591
|
Mon Sep 21 15:57:08 2020 |
Jordan | Update | VAC | TP3 Forepump Replacement and Vac reset |
I removed the forepump (Varian SH-110) for TP3 today to see why it had failed over the weekend. I tested it in the C&B lab and the ultimate pressure was only ~40torr. I checked the tip seals and they were destroyed. The scroll housing also easily pulled off of the motor drive shaft, which is indicative of bad bearings. The excess travel in the bearings likely led to significant increase in tip seal wear. This pump will need to be scrapped, or rebuilt.
I tested the spare Varian SH-110 pump located at the X-end and the ultimate pressure was ~98 mtorr. This pump had tip seals replaced on 11/5/18, and is currently at 55163 operating hours. It has been installed as the TP3 forepump.
Once installed, restarting the pump line occured as follows: V5 Closed, VA6 closed, VASE Closed, VASV closed, VABSSCI closed, VABS closed, VABSSCO closed, VAEV closed, VAEE closed,TP3 was restarted and once at normal operation, valves were opened in same order.
The pressure differential interlock condition for V5 was temporaily changed to 10 torr (by Gautam), so that valves could be opened in a controlled manner. Once, the vacuum system was back to normal state the V5 interlock condition was set back to the nominal 1 torr. Vacuum system is now running normally. |
Attachment 1: Screenshot_from_2020-09-21_15-56-04.png
|
|
Attachment 2: 20200921_145043.jpg
|
|
Attachment 3: 20200921_145040.jpg
|
|